Solve 3x ( x+3) = 2 (5x+1).
Answer by KnightOwlTutor(293)

Multiply out each side of the equation
3x^2+9x=10x+2
subtract 10x from both sides
3x^2-x=2
subtract 2 from both sides
3x^2-x-2=0
Factor the quadratic equation.

(3x+1)(x-2)=0
x=-1/3 or x=2